40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 78     Entry time: Thu Feb 25 19:03:59 2010
Author: Tara Chalermsongsak 
Type: Laser 
Category: RC noise 
Subject: Noise floor from PD and SR560 

Noise floor from PD and SR560 are measured, then converted from V/rHz to frequncy/rHz by the calibration from IFR2023 (0.7 MHz/Volt).

 SR560 has the same setting (gain invert 100, low pass filter at 1 Hz) as it did yesterday (when I measured the power spectrum

of the feed back signal.) The power spectrum of the feedback signal (from yesterday) is plotted in gray,  noise floor from PD( mixed with 160MHz signal from IFR2023) is plotted in pink, and noise from SR560 is in blue. I'll get the RC noise data on 2010Feb09, so I can plot them on the same graph.

Attachment 1: noise_floor.png  12 kB  | Hide | Hide all
noise_floor.png
ELOG V3.1.3-